OpenStudy (anonymous):

Which of the following is the slope between the two points (5, 1) and (4, 6)? one fifth -one fifth 5 -5

OpenStudy (anonymous):

slope=y2-y1/x2-x1=6-1/4-5 =5/-1 =-5

OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 2x - y = 8. intercept = (2, 0); y-intercept = (0, -8) x-intercept = (4, 0); y-intercept = (0, -8) intercept = (2, 0); y-intercept = (0, 8) intercept = (4, 0); y-intercept = (0, 8)